OBSERVATION AND ANALYSIS OF THE SLIT JET/FTIR SPECTRA OF SOME SMALL DIMERS

\underline{ROBERT L. SAMS} , THOMAS A. BLAKE, STEVEN W. SHARPE, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, 3020 Q Avenue, Richland, WA 99352 (PNNL is operated for the US Department of Energy by the Battelle Memorial Institute under contract DE-AC06-76RLO 1830).

As a demonstration of our cw slit jet/FTIR spectrometer we have recorded and analyzed the jet (T_{rot} = 10 K) spectra of N_{2}O dimer, N_{2}O-Ar dimer, CO_{2} dimer, and CO_{2}-Ar dimer in the 2100 to 2400 cm^{−1}region. A Bruker IFS 120HR FTIR spectrometer is used to interrogate the molecular and cluster species in a cw jet formed by the expansion of argon and CO_{2} or N_{2}O through a 12 cm × 50 μm slit. Preliminary analysis of the N_{2}O dimer spectrum gives the following constants: A′′ = 0.299421(5) cm^{−1}, B′′ = 0.059928(1) cm^{−1}, C′′ =0.049841(1) cm^{−1}, Δ′′_{J} = 0.192(3) × 10^{−6} cm^{−1}, Δ′′_{JK} = −0.93(3) × 10^{−6} cm^{−1}, Δ′′_{K} = 0.71(1) × 10^{−5} cm^{−1}, δ′′_{J} = 0.30(9) × 10^{−7} cm^{−1}, δ′′_{K} = 0.7(2) × 10^{−6} cm^{−1}and for the upper state A′ = 0.299067(4) cm^{−1}, B′ = 0.059665(1) cm^{−1}, C′ = 0.049647(1) cm^{−1}, Δ′_{J} = 0.195(3) × 10^{−6} cm^{−1}, Δ′_{JK} = −0.96(2) × 10^{−6} cm^{−1}, Δ′_{K} = 0.712(8) × 10^{−5} cm^{−1}, δ′_{J} = 0.33(9) × 10^{−7} cm^{−1}, δ′_{K} = 0.9(1) × 10^{−6} cm^{−1}, and ν_{0} = 2229.48295(2) cm^{−1}. Spectroscopic constants for the other dimers will also be presented.